# Sample size and power calculations for causal inference in observational studies

### v3

@@ 3.5 Analytical formula for sample size calculation
Summarizing the building blocks in Section 3.2-3.4, we propose to approximate the distribution of the propensity score and the potential outcomes as
$e(X)=\mathrm{expit}(W_{e}),\quad Y(z)=a_{z}W_{e}+\epsilon_{z},\quad W_{e}\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2}),\quad\epsilon_{z}\sim\mathsf{N}(\mu_{z},\sigma_{z}^{2}),$ (16)
respectively, where the propensity score model parameters $(\mu_{e},\sigma_{e})$ and the outcome model parameters $(a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2})$ can be uniquely determined by the user specified inputs $(r,\phi)$ and $(r,\phi,E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1}^{2},\rho_{0}^{2})$, respectively. To further reduce the number of inputs, we impose $S_{1}^{2}=S_{0}^{2}=S^{2}$ and $\rho_{1}=\rho_{0}=\rho$. Empirical evidence from multiple real data, e.g. the ones investigated in the supplementary materials, supported this restriction. Furthermore, following the convention in previous power analysis, we define the standardized effect size as the treatment effect scaled by the standard error: $\widetilde{\tau}=\tau/S$.
The following theorem presents the analytical formula for the sample size calculation of the $\tau^{\mbox{\tiny{ATE}}}$ (the proof is relegated to Supplementary Material A.6).
@@
Let $\widetilde{\tau}$ be the desired standardized effect size of the ATE. The minimum sample size for a two-sided level-$\alpha$ test (4) for detecting $\widetilde{\tau}$ with power $\beta$ is $N=\widetilde{V}\left(z_{1-\alpha/2}+z_{\beta}\right)^{2}/{\widetilde{\tau}^{2}}$, with
$\widetilde{V}=$ $\rho^{2}\left(\frac{\sigma_{e}^{2}}{\sigma_{e\mid 1}^{2}}+\frac{\sigma_{e}^{2}}{\sigma_{e\mid 0}^{2}}\right)+2(1-\rho^{2})+\left\{\rho^{2}\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid 1}^{2}}+(1-\rho^{2})\right\}\exp(-\mu_{e}+\sigma_{e}^{2}/2)$
+$+\left\{\rho^{2}\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid 0}^{2}}+(1-\rho^{2})\right\}\exp(\mu_{e}+\sigma_{e}^{2}/2),$ (17)
where $\sigma_{e\mid z}^{2}=\mathbb{V}(W_{e}\mid Z=z)$.
@@
Theorem 2 shows that the formula of $\widetilde{V}$ does not depend on $\mu_{z}$. Moreover, equation (12) shows that $E_{z}$ is involved in solving $\mu_{z}$ but not $a_{z}$ and $\sigma_{z}^{2}$. Therefore, calculating $\widetilde{V}$ does not require $E_{z}$, further reducing the number of inputs. Indeed, the formula of $\widetilde{V}$ only involves $\sigma_{e}^{2}$, $\mu_{e}$ and $\sigma_{e\mid z}^{2}$, all of which can be derived from the distribution of $e(X)$ that is determined by $(r,\phi)$, and an additional input $\rho$. This simplification reduces the number of inputs for the sample size calculation to two, namely $(\phi,\rho)$, in addition to the standard $\widetilde{\tau},\alpha,\beta,r$. In the special case of a randomized trial, $\rho=0$, and the minimal sample size reduces to formula (6).
@@
Remark 1. The above analytical derivation elucidates the connection and distinction between sample size calculations in randomized trials and observational studies. Specifically, four parameters — $\alpha$, $\beta$, the standardized effect size $\widetilde{\tau}$, and the treatment proportion $r$—are required in both cases. For observational studies, two additional inputs are required: the overlap coefficient $\phi$ and the correlation parameter $\rho^{2}$, both of which have natural interpretation. This result is intuitive because the key difference between observational studies and randomized trials is the presence of confounders in the former, which are pre-treatment covariates that affect the treatment assignment and the outcome simultaneously. The parameters $\phi$ and $\rho^{2}$ quantify the strength of confounder-treatment and confounder-outcome correlation, respectively. Similar strategy is used in the formulation of the E-value (Ding and VanderWeele,, 2016). Intuitively, as the overlap decreases (i.e., $\phi$ deceases) and the correlation between covariates and the outcome increases (i.e., $\rho^{2}$ increases), the observational study deviates further from a randomized trial, rendering the marginal distribution of the observed outcomes deviate further from that of the potential outcomes and thus require more sample units to recover the causal estimands. Indeed, the calculated sample size monotonically increases with $\rho^{2}$ from (17) (the proof is relegated to Supplementary Material A.7.). Moreover, numerical experiments show that the calculated sample size monotonically decreases with $\phi$.
@@
Remark 2. The sample size formula of Shook-Sa and Hudgens, (2022) – derived based on the variance inflation factor – is a special case of formula (17). Specifically, when assuming the propensity score distribution $e(X)\sim\operatorname{\mathrm{e}xpit}(\mathsf{N}(\mu_{e},\sigma_{e}^{2}))$, their formula becomes $\widetilde{V}_{\mathrm{SH}}=2+\exp(-\mu_{e}+\sigma_{e}^{2}/2)+\exp(\mu_{e}+\sigma_{e}^{2}/2)$, which is formula (17) with $\rho=0$. We show that variance $\widetilde{V}_{\mathrm{SH}}$ is smaller than the variance $\widetilde{V}$ in (17) by a term proportional to $\rho^{2}$ (details in Supplementary Material A.7). Although $\rho^{2}$ is small in many of our simulation studies, ignoring it risks underestimating the sample size in general settings. Therefore, we recommend to report power calculations with a range of $\rho^{2}$ values between 0 and $R^{2}_{Y|X}$. Shook-Sa and Hudgens, (2022) discussed a remainder term that implicitly adjusts for the confounder-treatment association, but did not point out the connection, and the remainder term is only estimable with prior data on both weights and outcomes. In contrast, our method does not require this information and explicitly formulates the role of the confounder-treatment association.
@@
Remark 3. For binary outcomes, because the mean of a Bernoulli distribution determines the variance, compared to the continuous outcomes in Section 3.3, we have two fewer constraints to determine the outcome distribution. As a simple solution, we recommend to use linear models for binary outcomes so that the proposed method directly applies. Numerical examples in Section 4.2 and Supplementary Material D.3 suggest that this approach generally works well in practice.
4 Numerical Examples
4.1 Synthetic Data
@@
Following the simulation design in Li et al., (2019), we simulate $N=1,000,000$ units with 10 covariates of different types as the superpopulation. Covariates $X_{1}$ to $X_{4}$ are binary with $\mathsf{Ber}(p)$ with $p=0.2,0.4,0.6,0.8$ respectively; $X_{5}\sim\mathsf{Unif}(0,1)$; $X_{6}$ to $X_{8}$ follows a Poisson distribution with mean parameters $1,2,$ and $3$ respectively; $X_{9}\sim\mathsf{Gamma}(2,3)$ and $X_{10}\sim\mathsf{Beta}(2,3)$. We generate the treatment indicator $Z\sim\mathsf{Ber}(\operatorname{\mathrm{e}xpit}(\beta_{0}+\kappa X\beta))$, where $\beta=(1,1,-1,0,-2,1,0.5,0,0,0)^{\prime}$ and $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$ is the parameter that controls the overlap between two treatment groups. For these $\alpha$ values, we choose $\beta_{0}=\{0,-0.248,-0.489,-0.722,-0.860,-0.951\}$ accordingly so that the treatment proportion $r=\mathbb{E}(Z)=0.5$. We then simulate the potential outcome $Y(z)\sim\mathsf{N}(X\gamma+\tau z,4^{2})$, where $\gamma=(1,1,-1,-1,0,-1,-1,0,1,1)^{\prime}$, and let the observed outcome $Y_{i}=Z_{i}Y(1)+(1-Z_{i})Y(0)$. The outcome model implies a homogeneous treatment effect $\tau(x)=1$ and thus all average effects including $\tau^{\mbox{\tiny{ATE}}}$ equal 1.
The overlap between two treatment groups is shown in Figure 2: $\kappa=0$ corresponds to randomized trials; as $\kappa$ increases, the overlap decreases. The values of $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$ correspond to the overlap coefficient $\phi\in\{1.00,0.98,0.93,0.87,0.84,0.81\}$, respectively.
@@
Figure 2: Distribution of the propensity scores in two treatment arms with fixed $r=0.5$ and various overlap coefficient $\phi=1.00,0.98,0.93,0.87,0.84,0.81$, respectively.
Now we fix the power at 0.80, and solve for the sample size under each value of $\kappa$. We calculate the eight summary statistics $(r,\phi,E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1}^{2},\rho_{0}^{2})$ as discussed in Section 3 from the $N=1,000,000$ units in the superpopulation. The proportion of treatment, $r$, is calculated by the sample mean $\overline{Z}$, and the observed mean and variance, $E_{1},E_{0},S_{1}^{2},S_{0}^{2}$ are calculated with values of $Y$ on subsets $Z=1,0$, respectively. With the logistic model for the propensity score, $e(X)=\operatorname{\mathrm{e}xpit}(\beta_{0}+X\beta)$, the latent variable $W_{e}$ is estimated by $\widehat{\beta}_{0}+X\widehat{\beta}$, where $(\widehat{\beta}_{0},\widehat{\beta})$ is the estimated intercept and coefficients of the logistic regression of $Z$ on $X$. Although none of the covariates $X$ is normal, the latent linear combination of $X$, $W_{e}$, appears to be close to normal (see the rightmost column of Figure 4 in Supplementary Material B.1), and the correlation parameters $\rho_{z}$ can be estimated by the sample correlation between estimated $W_{e}$ and $Y(z)$. We also calculate the pooled variance $S^{2}$ and pooled correlation $\rho$ empirically. To calculate $\phi$, we use the empirical cumulative probability function of the propensity scores (details are relegated to Supplementary Material B.3). We present $(r,\phi)$ in Table 1 below, and the remaining quantities $(E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1}^{2},\rho_{0}^{2})$ and the pooled quantities $(S^{2},\rho^{2})$ are shown in Table 2 in Supplementary Material C.2. We also calculate and show in Table 2 the R-squared statistic $R^{2}$ of linear model $Y\sim X$, with $Y$ centered within the respective treatment arm. The results show that $S_{1}^{2}\approx S_{0}^{2}\approx S^{2}$ and $\rho_{1}^{2}\approx\rho_{0}^{2}\approx\rho^{2}$ in all settings. The $R^{2}$ turns out to be much larger than $\rho^{2}$, which is not surprising as $R^{2}$ is the theoretical upper bound of $\rho^{2}$, and provides the most conservative estimate when there is no substantive knowledge on the range of $\rho^{2}$.
@@
$\phi$ Size True PS Est PS Size True PS Est PS Size True PS Est PS
1.00 645 0.83 0.88 629 0.80 0.87 629 0.80 0.87
0.98 656 0.79 0.87 628 0.78 0.85 653 0.80 0.88
0.93 754 0.79 0.87 626 0.73 0.80 742 0.78 0.86
0.88 993 0.78 0.85 624 0.60 0.65 950 0.77 0.84
0.84 1290 0.80 0.85 623 0.48 0.53 1216 0.77 0.84
0.81 1617 0.81 0.87 622 0.40 0.45 1481 0.79 0.81
Table 1: Sample size calculation by the proposed method, the two-sample $z$-test and the method in (Shook-Sa and Hudgens,, 2022) to achieve nominal power of 0.80 under various degrees of overlap. The power is obtained from random sampling of the calculated size from the $N=1,000,000$ units. The Monte Carlo error for a power of 0.8 is 0.004.
@@
We use inputs $(r,\phi,\rho^{2})$ and standardized effect size $\widetilde{\tau}=1/S$ to inversely calculate the sample size $N_{\mathrm{pos}}$. Table 1 shows the calculated sample size to achieve the nominal power 0.80 under various degrees of overlap. Then, we sample $N_{\mathrm{pos}}$ units randomly from the $N=1,000,000$ units in the superpopulation. We repeat the process for $B=10,000$ times; each time, we conduct the hypothesis test (4) at level $\alpha=0.05$, with $\widehat{\tau}_{N}$ calculated with true and estimated propensity scores, respectively. We then estimate the power by the proportion of the $B=10,000$ replications where the null hypothesis is rejected. Table 1 also reports the power calculated based on Shook-Sa and Hudgens, (2022). The proposed method consistently leads to powers close to the nominal 0.80; the method of Shook-Sa and Hudgens, (2022) leads to similar power. As noted before, the power based on the estimated propensity scores is higher than that based on the true scores, rendering the proposed method conservative. Note that the discrepancy in variance between the true and estimated propensity scores could be magnified in the simulation compared to most real studies, because both $Z$ and $Y$ are drawn from the underlying true model. We also include the sample size calculated by two-sample $z$-test for comparison. When the overlap is perfect, the $z$-test correctly calculates the sample size. However, when the overlap is poor, meaning that the intended observational study is far from a target randomized trial, the $z$-test, which treats observational studies as randomized trials and thus ignore the degree of overlap, tends to vastly underestimate the sample size.
@@
The parameter $\rho^{2}$ is usually unknown; we suggest treating $\rho^{2}$ as a sensitivity parameter between 00 and $R^{2}$. For $\phi\in\{1.00,0.98,0.93,0.88,0.84,0.81\}$, sensitivity analyses will yield sample sizes of $(629,630)$, $(653,669)$, $(742,821)$, $(959,1239)$, $(1225,1818)$ and $(1515,2516)$, respectively. The results show that when the overlap is good, i.e., the observational study resembles a randomized trial, the minimal sample size is robust to the choice of $\rho^{2}$. However, as the overlap deteriorates, the sample size becomes more sensitive to the choice of $\rho^{2}$. This finding restates the importance of covariate overlap in observational studies, and presents the challenge of sample size calculation with worse overlap. The wide intervals do not reflect a drawback of the proposed method; instead, it displays the sensitivity to the “hidden” parameter $\rho^{2}$, which has been mostly ignored in the existing literature. The sensitivity also shows the importance of the bound $\rho^{2}\leq R^{2}$, because $R^{2}$ is often small (e.g. 0.2) in clinical studies, which guarantees that the sample size falls into a reasonable range with modest overlap.

### v4

@@ 3.4 Analytical formula for sample size calculation
In this subsection, we will build on Section 3.2-3.3 to derive an analytical sample size calculation formula. Following the convention of power analysis, we define the effect size as the treatment effect standardized by the standard error: $\widetilde{\tau}=\tau/S$. The following theorem presents the analytical formula for the sample size calculation of $\tau^{\mbox{\tiny{ATE}}}$ (the proof is relegated to Appendix A.7).
@@
Let $\widetilde{\tau}$ be the desired standardized effect size of the ATE. The minimum sample size for the one-sided level-$\alpha$ test defined in (4) for detecting $\widetilde{\tau}$ with power $\beta$ is $N=\widetilde{V}\left(z_{1-\alpha}+z_{\beta}\right)^{2}/{\widetilde{\tau}^{2}}$, with
$\widetilde{V}=2\left\{1+(\rho^{2}\sigma_{e}^{2}+1)\exp\left(\frac{\sigma_{e}^{2}}{2}\right)\cosh(\mu_{e})\right\},$ (14)
where $\cosh(\mu_{e})=\{\exp(\mu_{e})+\exp(-\mu_{e})\}/2$.
@@
The minimum sample size for the two-sided counterpart of test (4) is obtained by replacing $z_{1-\alpha}$ with $z_{1-\alpha/2}$ in Theorem 2. Equation (14) shows that the variance $\widetilde{V}$ only involves $\mu_{e}$ and $\sigma_{e}^{2}$, both of which can be derived from the distribution of $e(X)$ that is determined by $(r,\phi)$, and $\rho$. This simplification reduces the number of inputs for the sample size calculation to two, namely $(\phi,\rho)$, in addition to the standard $\widetilde{\tau},\alpha,\beta,r$. In the special case of a randomized trial, $\rho=0$, and the minimal sample size reduces to formula (6). Based on Equation (14) and jointly with Proposition 1(ii) and the relation in (8), the following corollary on the relationship between $N$ and $(\rho,\phi)$ is immediate.
Corollary 1.
The variance $\widetilde{V}$ and thus the sample size $N$ (i) monotonically increases in $|\rho|\in(0,1]$ given $(r,\phi)$, and (ii) monotonically decreases in $\phi\in(0,1]$ given $r$.
Figure 2 shows a flow chat of how to decompose the variance of the Hájek estimator into three components and compute each component from the proposed inputs.
[…]
Figure 2: Flow chat of decomposing and computing the variance of the Hájek estimator from summary inputs.
Remark 1. In summary, our derivation relies on (i) a logistic model for the propensity scores, and (ii) a semiparametric homoscedastic restricted mean outcome model with a homogeneous treatment effect. In real applications, the true data generating process might deviate from the assumptions. In particular, heterogeneous treatment effects are common. We can extend the homogeneous outcome model (11) to a varying coefficient model to accommodate heterogeneity: $Y(z)=f(X)+z\tau(X)+\epsilon$, where $\tau(X)$ stands for the conditional average treatment effect at covariate $X$. Maintain the homoscedasticity and joint normality assumptions and use the same projection to the space spanned by $W_{e}$ as in the homogeneity model. Then, simple algebra (details are omitted) shows that the asymptotic variance of the Hájek estimator $V$ of the heterogeneity model depends on the variance of $\tau(X)$, the knowledge of which is generally not available before data collection as in power analysis. This challenge is also discussed in the context of sample size calculation for rerandomization (Branson et al., 2024). Instead, we use simulations in Section 5 to examine consequences of violation to homogeneity and find Formula (14) is generally robust.
@@
Remark 2. Our analytical derivation elucidates the connection and distinction between power calculations in randomized trials and observational studies. Specifically, both cases require four parameters: $\alpha$, $\beta$, the standardized effect size $\widetilde{\tau}$, and the treatment proportion $r$. Observational studies require two additional inputs: the overlap coefficient $\phi$ and the correlation parameter $\rho^{2}$, both of which have a natural interpretation. This result is intuitive because the key difference between observational studies and randomized trials is the presence of confounders in the former, which are pre-treatment covariates that affect the treatment assignment and the outcome simultaneously. The parameters $\phi$ and $\rho^{2}$ quantify the strength of confounder-treatment and confounder-outcome association, respectively. A similar strategy is used in the formulation of the E-value (Ding and VanderWeele, 2016). Intuitively, as the covariate overlap between groups (encoded by $\phi$) decreases and the correlation between covariates and the outcome (encoded by $\rho^{2}$) increases, an observational study deviates further from a randomized trial, rendering the marginal distribution of the observed outcomes deviate further from that of the potential outcomes and thus require more sample units to identify the causal estimands. This matches Corollary 1.
@@
Remark 3. The sample size formula in Shook-Sa and Hudgens (2022), derived based on a variance inflation factor, is a special case of formula (14). Specifically, assuming $e(X)\sim\operatorname{\mathrm{e}xpit}(\mathsf{N}(\mu_{e},\sigma_{e}^{2}))$, their formula becomes $\widetilde{V}_{\mathrm{SH}}=2\left\{1+\exp(\sigma_{e}^{2}/2)\cosh(\mu_{e})\right\}$. Comparing to formula (14), we have $\widetilde{V}=\widetilde{V}_{\mathrm{SH}}+2\rho^{2}\sigma_{e}^{2}\exp(\sigma_{e}^{2}/2)\cosh(\mu_{e})$. Therefore, $\widetilde{V}_{\mathrm{SH}}$ equals $\widetilde{V}$ when $\rho=0$ and is smaller than $\widetilde{V}$ otherwise. Although $\rho^{2}$ is small in many of our simulation studies, ignoring it risks underestimating the sample size in general settings. Shook-Sa and Hudgens (2022) discussed a term that implicitly adjusts for the confounder-treatment association, but it is not estimable without prior data on both weights and outcomes.
@@
Remark 4. For binary outcomes, because the mean of a Bernoulli distribution determines the variance, compared to continuous outcomes in Section 3.3, we have two fewer constraints to determine the outcome distribution. As a simple solution, we recommend using linear models for binary outcomes so that the proposed method directly applies. Numerical examples in Section 5.3 and Appendix C.1 suggest that this approach generally works well in practice.
4 Weighted average treatment effect estimands
In the presence of poor overlap between the treatment groups, the inverse probability weighting can lead to excessive variance due to extreme propensity scores. Alternative weighting schemes are often more desirable. In this section, we extend the above methodology to a general class of weighted average treatment effect (WATE). Assume that the observed sample is drawn from a population with covariates distribution $f(x)$, and let $f(x)h(x)$ denote the covariate distribution of a target population, where $h(x)$ is called a tilting function. Then, we can represent the average treatment effect on that target population using a WATE
$\tau^{h}=\frac{\operatorname{\mathbb{E}}[\tau(x)h(x)]}{\operatorname{\mathbb{E}}[h(x)]}.$ (15)
When $h(x)=1$, $\tau^{h}$ reduces to ATE; when $h(x)=e(x)$, $\tau^{h}$ becomes the average treatment effect for the treated (ATT); when h(x)=e(x)(1e(x))h(x)=e(x)(1-e(x)), $\tau^{h}$ becomes the average treatment effect for the overlap population (ATO). We can use a unified weighting strategy to identify the WATE because $\tau^{h}=\operatorname{\mathbb{E}}[w_{1}(X)ZY-w_{0}(X)(1-Z)Y]$, where $w_{1}(x)=h(x)/e(x),w_{0}(x)=h(x)/(1-e(x))$ (Li et al., 2018a). This corresponds to (i) the inverse probability weight (IPW), $(w_{1}=1/e(x),w_{0}=1/(1-e(x)))$, for the ATE; (ii) the ATT weight, $(w_{1}=1,w_{0}=e(x)/(1-e(x)))$, for the ATT; and (iii) the overlap weight (OW), $(w_{1}=1-e(x),w_{0}=e(x))$, for the ATO.
The Hájek estimator for the WATE is:
$\widehat{\tau}_{w}=\frac{\sum_{i}w_{1}(X_{i})Z_{i}Y_{i}}{\sum_{i}w_{1}(X_{i})Z_{i}}-\frac{\sum_{i}w_{0}(X_{i})(1-Z_{i})Y_{i}}{\sum_{i}w_{0}(X_{i})(1-Z_{i})}.$ (16)
We can adapt the procedures in Section 3 to derive sample size calculation formulas for the ATT and ATO estimands. Specifically, the procedures to determine the propensity score and the outcome distributions remain the same, but the variance formula changes. Unlike the ATE, there is no deterministic order between the variance of the Hájek estimator with the estimated versus true propensity score for other WATE estimands (Yang and Ding, 2018). For the same reasons as with the ATE stated in Section 2, we choose to proceed with the variance with the true $e$.
We show the variance of the Hájek estimator for the WATE with the tilting function $h$ is:
$V_{w}\equiv\operatorname{\mathbb{V}}(\widehat{\tau}_{w})=\frac{1}{\mathbb{E}[h(X_{i})^{2}]}\mathbb{E}\left[\left\{\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}+\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}\right\}h(X_{i})^{2}\right],$ (17)
where $\xi_{z}=\mathbb{E}_{h}[Y_{i}(z)]=\mathbb{E}[h(X_{i})Y_{i}(z)]/\mathbb{E}[h(X_{i})]$; the proof is relegated to Appendix A.8. For the ATE, $h(X_{i})=1$, and $V_{w}$ reduces to $V$ as in Equation 3. For the ATT and ATO, $h(X_{i})$ is a function of $e(X_{i})$ and thus $V_{w}$ is also fully determined by the joint distribution of $e(X_{i})$ and the potential outcomes $Y_{i}(z)$, which can be computed as in Sections 3.2 and 3.3. However, neither the ATT or the ATO has a closed-form expression of variance, but they can be computed numerically.
We outline how to compute $V_{w}$ with the tilting function $h(X)$ given the parameters $(\mu_{e},\sigma_{e})$ and $(a,\sigma_{y}^{2})$. For the ATT and ATO, $h(X)$ is a function of the propensity score $e(X)$, whose distribution can be approximated by $W_{e}$. Therefore, we can also write the tilting function as $h(W_{e})=\operatorname{\mathrm{e}xpit}(W_{e})$ for ATT and $h(W_{e})=\operatorname{\mathrm{e}xpit}(W_{e})[1-\operatorname{\mathrm{e}xpit}(W_{e})]$ for ATO, such that $h(W_{e})$ has approximately the same distribution as $h(X)$. Calculating the denominator $\mathbb{E}[h(X)^{2}]=\mathbb{E}[h(W_{e})^{2}]$ where $W_{e}\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$ involves numerically (e.g. Simpson’s method) evaluating the following integral
$\int_{-\infty}^{\infty}h(u)^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u,$ (18)
$\mathbb{E}\left[\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}h(X_{i})^{2}\right]$ (19)
$=$ $a^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{y}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right],$
$\mathbb{E}\left[\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}h(X_{i})^{2}\right]$ (20)
$=$ $a^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{y}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right],$
5 Numerical Experiments
5.1 Synthetic Data: homogeneous treatment effect
@@
Following the simulation design in Li et al. (2019), we simulate $N=1,000,000$ units with 10 covariates of different types as the superpopulation. Covariates $X_{1}$ to $X_{4}$ are binary with $\mathsf{Ber}(p)$ with $p=0.2,0.4,0.6,0.8$, respectively; $X_{5}\sim\mathsf{Unif}(0,1)$; $X_{6}$ to $X_{8}$ follows a Poisson distribution with mean parameters $1,2,$ and $3$ respectively; $X_{9}\sim\mathsf{Gamma}(2,3)$ and $X_{10}\sim\mathsf{Beta}(2,3)$. We generate the treatment indicator $Z\sim\mathsf{Ber}(\operatorname{\mathrm{e}xpit}(\beta_{0}+\kappa X\beta))$, where $\beta=(1,1,-1,0,-2,1,0.5,0,0,0)^{\prime}$. The parameter $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$ controls the overlap between two treatment groups, corresponding to the overlap coefficient $\phi\in\{1.00,0.98,0.93,0.87,0.84,0.81\}$, respectively; $\kappa=0$ corresponds to a randomized trial. Figure 3 shows the overlap between two groups with varying $\kappa$. For these $\kappa$ values, we choose $\beta_{0}\in\{0,-0.248,-0.489,-0.722,-0.860,-0.951\}$ accordingly so that the treatment proportion $r=\mathbb{E}(Z)=0.5$. We simulate from a potential outcome model with a homogeneous treatment effect $\tau$: $Y(z)\sim\mathsf{N}(X\gamma+\tau z,4^{2})$, where $\gamma=(1,1,-1,-1,0,-1,-1,0,1,1)^{\prime}$ and $\tau=1$, and let the observed outcome $Y_{i}=Z_{i}Y(1)+(1-Z_{i})Y(0)$.
@@
Figure 3: Distribution of the propensity scores in two treatment arms with fixed $r=0.5$ and various overlap coefficient $\phi=1.00,0.98,0.93,0.87,0.84,0.81$, respectively, in Simulation 5.1.
[…]
Table 1: Sample size calculation by the proposed method, the two-sample $z$-test and the method in (Shook-Sa and Hudgens, 2022) to achieve nominal power of 0.80 under various degrees of overlap. The power is obtained from random sampling of the calculated size from the $N=1,000,000$ units. The Monte Carlo error for a power of 0.8 is 0.004. The treatment proportion $r$ is fixed at 0.5. The empirical $S^{2}$ is close to 2020 in all settings.
@@
$\phi$ $\rho^{2}$ $R^{2}$ Size True PS Est PS Size True PS Est PS Size True PS Est PS
1.00 0.02 0.20 645 0.83 0.88 629 0.80 0.87 629 0.80 0.87
0.98 0.03 0.20 656 0.79 0.87 628 0.78 0.85 653 0.80 0.88
0.93 0.03 0.19 754 0.79 0.87 626 0.73 0.80 742 0.78 0.86
0.88 0.02 0.19 993 0.78 0.85 624 0.60 0.65 950 0.77 0.84
0.84 0.02 0.19 1290 0.80 0.85 623 0.48 0.53 1216 0.77 0.84
0.81 0.02 0.19 1617 0.81 0.87 622 0.40 0.45 1481 0.79 0.81
@@
We use inputs $(r,\phi,\rho^{2})$ and the standardized effect size $\widetilde{\tau}=1/S$ to inversely calculate the sample size $N_{\mathrm{pos}}$. Table 1 shows the calculated sample size to achieve the nominal power 0.80 under various degrees of overlap. Then, we sample $N_{\mathrm{pos}}$ units randomly from the $N=1,000,000$ units in the superpopulation. We repeat the process for $B=10,000$ times; each time, we conduct the hypothesis test (4) at level $\alpha=0.05$, with $\widehat{\tau}_{N}$ calculated with true and estimated propensity scores, respectively. We then estimate the power by the proportion of the $B=10,000$ replications where the null hypothesis is rejected. Table 1 also reports the power calculated based on Shook-Sa and Hudgens (2022) (SH). The proposed method consistently leads to powers close to the nominal 0.80; the SH method leads to slightly lower power. The similarity between the results from our method and from the SH method is as expected from Remark 3 given the small $\rho^{2}$. As noted before, the power based on the estimated propensity scores is higher than that based on the true scores, rendering the proposed method conservative. Note that the discrepancy in variance between the true and estimated propensity scores could be magnified in the simulation compared to most real studies, because both $Z$ and $Y$ are drawn from the underlying true model. We also include the sample size calculated by two-sample $z$-test for comparison. When the overlap is perfect, the $z$-test correctly calculates the sample size. However, when the overlap is poor, meaning that the intended observational study is far from a target randomized trial, the $z$-test, which treats observational studies as randomized trials and thus ignores the degree of overlap, tends to vastly underestimate the sample size.
@@
The parameter $\rho^{2}$ is usually unknown; we suggest treating $\rho^{2}$ as a sensitivity parameter between 0 and $R^{2}$. For $\phi\in\{1.00,0.98,0.93,0.88,0.84,0.81\}$, sensitivity analyses will yield sample sizes between $(629,630)$, $(653,669)$, $(742,821)$, $(959,1239)$, $(1225,1818)$ and $(1515,2516)$, respectively. The results show that when the overlap is good, i.e., the observational study resembles a randomized trial, the minimal sample size is robust to the choice of $\rho^{2}$. However, as the overlap deteriorates, the sample size becomes more sensitive to the choice of $\rho^{2}$. This finding restates the importance of covariate overlap in observational studies and presents the challenge of sample size calculation with worse overlap. The wide intervals do not reflect a drawback of the proposed method; instead, it displays the sensitivity to the “hidden” parameter $\rho^{2}$, which has been mostly ignored in the existing literature. Such sensitivity also highlights the importance of the bound $\rho^{2}\leq R^{2}$, which guarantees that the sample size falls into a reasonable range with modest overlap because $R^{2}$ is often small (e.g. 0.2) in clinical studies.